OpenStudy (anonymous):

A tugboat goes upstream 160 miles in 10 hours. The return trip downstream takes 5 hours. Find the speed of the tugboat without a current and the speed of the current.

OpenStudy (anonymous):

Thank you

hero (hero):

Let me know what you get for b and c

OpenStudy (anonymous):

I got 16 for b and 32 for c

hero (hero):

How is it possible that b < c when we need b - c to be positive?

OpenStudy (anonymous):

I am not sure. honestly I have been stuck on this question for almost an hour.

hero (hero):

Hint: use d = rt, where r = (b ± c) 160 = 10(b-c) 160 = 5(b+c) Solve the system

hero (hero):

Try that and see what you get

OpenStudy (anonymous):

I got c=16 and b=32. so the tugboat goes 32 mph and the speed of the current would be 16. Thank you for your help.

hero (hero):

No, sorry, that still is not correct. I have no clue how you are going about solving this

hero (hero):

I show you that 160 = 10(b-c) 160 = 5(b+c) The next step would be to divide both sides of the first equation by 10, and both sides of the second equation by 5 to get: 16 = b - c 32 = b + c However, we need to find b and c separately.

OpenStudy (anonymous):

I am solving it by eliminating the c or b. Like I stated above, I have been at this equation for an hour just about, I am confused on how to solve for this. I am sorry for miss understanding you, I just do not get it.

hero (hero):

I'm pretty sure you got 16 = b - c 32 = b + c in your initial steps. The problem is, we are not done yet.

hero (hero):

We're not done until we isolate one of the variables and find a specific value for it

hero (hero):

If you think about it, we have a system of equations here. And we can easily eliminate c by adding both equations together. If we do that we get 48 = 2b

OpenStudy (anonymous):

so if 48=2b, we would divide both sides by 2?

hero (hero):

Yes, but the important thing is that you understand how I got to 48 = 2b

OpenStudy (anonymous):

you took 16=b-c 32=b+c and you added both sides together, so that is how you got 48=2b. 16+32=48 and b+b 1+1= 2 I have noticed you could not do this to c, because a negative plus a positive would eliminate the c all together.

hero (hero):

Okay, so have you actually solved for b yet?

hero (hero):

If so, what did you get?

OpenStudy (anonymous):

b=24

hero (hero):

Okay, so if b = 24, what is c?

OpenStudy (anonymous):

c would equal 0 if we used this term by adding the equations together.

hero (hero):

That was only to eliminate c so that we could solve for b. That doesn't mean c equals zero.

hero (hero):

c could equal any number. The way you get zero is by taking whatever number you have and subtracting that same number to get zero. 10 - 10 = 0 16 - 16 = 0 4 - 4 =0 a - a = 0 b - b = 0 those are just examples, but they show that those are numbers that are subtracted to get zero. But of course 10 ≠ 0, 16 ≠ 0, 4 ≠ 0, a ≠ 0, b ≠ 0, and in this case c ≠ 0

OpenStudy (anonymous):

would c=8 if we plugged 24 in an equation?

hero (hero):

Actually, it is 8. My bad. Good job

hero (hero):

Finally...you had me confused at certain points too

OpenStudy (anonymous):

Thank you, you had me re think myself, because both sides gave me 8

OpenStudy (anonymous):

so 24 would be the mph and 8 would be the current?

hero (hero):

24 mph = speed of the boat 8 mph = speed of the current That's why I used b and c as variables of course

OpenStudy (anonymous):

thank you for you help, I am glad we used it this way. you made it less confusing. :)

hero (hero):

As a general rule the speed of the boat must be greater than the speed of the current. Otherwise, the boat would not get to its destination.

OpenStudy (anonymous):

Okay. I will remember that.

hero (hero):

Also another thing you need to remember is this: Upstream = boat against current = b - c downstream = both with current = b + c This is why the equations are setup as such Upstream: 160 = 10(b-c) Downstream: 160 = 5(b+c)